OpenStudy (anonymous):

find a counterexample for the following statement: x^2 +x+41 is prime answers: x=1,x=3,x=5,x=7

OpenStudy (anonymous):

a counter example is the one that dosnt work when you plug it in...

OpenStudy (anonymous):

when I plug them in they all work 1^2+1+41=43 3^2+3+41=53 5^2 +5+41=71 7^2+7+41=97

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

which one of the answers is NOT prime?

OpenStudy (anonymous):

all of them are prime

OpenStudy (anonymous):

i think that questions wrong..

OpenStudy (anonymous):

yeah